Compute for the distance between the theodolite and the subtense bar if the subtended angle observed is 1degree 24 minutes and 32 seconds.
1
Expert's answer
2021-08-31T01:42:33-0400

Distance = "H= \frac{D}{2} \cot \frac{\alpha}{2}"

"\frac{\alpha}{2}= \frac{1^0+(\frac{24}{60})^0+(\frac{32}{60*60})^0}{2}=\frac{\frac{317}{225}}{2}=0.70444"

If D = 2 m

"H= \frac{2}{2} \cot 0.70444= 1.17659 m"